# Learn With Imagination: Safe Set Guided State-wise Constrained Policy Optimization

### v3

@@ 8 Conclusion and Future Prospectus
In this study, we introduce Safe Set Guided State-wise Constrained Policy Optimization (S-3PO), a novel algorithm pioneering state-wise safe optimal policies. This distinction is underlined by the absence of training violations, signifying an error-free learning paradigm. S-3PO employs a safeguard anchored in black-box dynamics to ensure secure exploration. Subsequently, it integrates a novel “imaginary” safety cost to guide the RL agent towards optimal safe policies. S-3PO outperforms existing methods in complex high-dimensional robotics tasks.
Nevertheless, a noteworthy limitation pertains to the potential costliness of acquiring a physical engine-based simulator (black-box dynamics model). A forward-looking perspective entails replacing the black-box dynamics model with a learned surrogate model, factoring in the nuances of errors in learned dynamics. This strategic move holds the promise of obliterating the final barrier impeding the seamless integration of safe RL training into real-world applications.
@@ A.1 Circumstances and Assumptions
Assumption 1 (2D Collision Avoidance).
@@ A.2 Safety Index Synthesis
Following the rules in (Liu & Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{A}_{S}^{D}(s)$ is always nonempty. By setting $\eta=0$, the parameterization rule of safety index design rule is defined as:
@@ B.1 S-3PO Algorithm
         if  $\mathbb{E}_{\hat{s}\sim\tau}[\mathcal{D}_{KL}(\pi_{\theta^{\prime}}\|\pi_{k})[\hat{s}]]\leq\delta$ and
$\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{\theta^{\prime}}}_{D}(\hat{s},a)-A^{\pi_{k}}_{D}(\hat{s},a)\right]\leq\mathrm{max}(-c,0)$ and
$(k\leq k_{\mathrm{safe}}~{}\mathrm{or}~{}\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{\theta^{\prime}}}(\hat{s},a)\right]\geq\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{k}}(\hat{s},a)\right])$ then
              $\theta_{k+1}\leftarrow\theta^{\prime}$triangleright Update policy
@@ B.2 Additional Algorithms
1:procedure AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\vec{v}^{r},S,S_{goal}$)
@@
1:procedure ISSA($\epsilon,\beta,\Sigma,n,a^{r}$)
2:     Phase 1: $\triangleright$ Phase 1
@@
20:end procedure
Appendix C Proof of Theorem 1
To prove the Theorem 1 we introduce two important corollaries to show that 1) the set of safe control is always nonempty if we choose a safety index that satisfies the design rule in (12); and 2) the proposed algorithm 3 is guaranteed to find a safe control if there exists one. With these two corollaries, it is then straightforward to prove the Theorem 1.
@@ C.1 Feasibility of the Safety Index for Continuous-Time System
Corollary 1 (Non-emptiness of the set of safe control).
@@
Note that the set of safe control $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$ is non-empty if and only if it is non-empty in the following two cases: $\phi(s)-\eta<0$ or $\phi(s)-\eta\geq 0$. In the following discussion, we first show that the safety index design rule guarantees a non-empty set of safe control if there’s only one obstacle when $\phi(s)-\eta\geq 0$ (Lemma 1). Then we show that the set of safe control is non-empty if there’s only one obstacle when $\phi(s)-\eta<0$ (Lemma 2). Finally, we leverage Lemma 1 and Lemma 2 to show $\mathcal{A}_{S}^{D}(s)$ is non-empty if there’re multiple obstacles at any state.
Lemma 1.
@@
The three cases cover all possible situations. Hence (17) always hold and the claim in the lemma is verified. ∎
Lemma 2.
@@
Now we show that $\dot{\phi}$ is bounded. According to the definition of safety index design rule, $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$. We ignored the subscript $i$ since it is assumed that there is only one obstacle. According to Assumption 1, we have the state space is bounded, thus both $d$ and $\dot{d}$ are bounded, which implies that $nd^{n-1}\dot{d}$ is bounded. Moreover, we have for all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$ according to Assumption 1, which indicates the mapping from $a$ to $(z,w)$ is surjective. Since $z$ and $w$ are bounded and both can achieve zeros according to Assumption 1, we have afor-$\forall a$, the corresponding $(z,w)$ are bounded. Since $\ddot{d}=-z\cos(\alpha)+v\sin(\alpha)w$, then $\ddot{d}$ is bounded. Hence $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ any $s$ that $\phi(s)-\eta<0$ and the claim is true. ∎
Proof of Corollary 1.
Proof.
If there is one obstacle, then lemma 1 and lemma 2 have proved that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for all $s$. Now we need to consider the case where there are more than one obstacle but the environment is sparse in the sense that at any time step, there is at most one obstacle which is safety critical, i.e. $\phi_{i}\geq 0$. To show nonemptiness of $\mathcal{A}_{S}^{D}(s)$, we consider the following two cases. In the following discussion, we set $\eta=0$ according to the safety index design rule.
@@
Denote ${\mathcal{A}_{S}^{D}}_{j}(s):=\{a\in\mathcal{A}\mid\phi_{j}(f(s,a))\leq\phi_{j}(s)\}$. Lemma 1 ensures that ${\mathcal{A}_{S}^{D}}_{j}(s)$ is nonempty. Denote ${\mathcal{A}_{S}^{D}}_{k}(s):=\{a\in\mathcal{A}\mid\phi_{k}(f(s,a))\leq 0\}$ where $k\neq j$. Since $\phi_{k}(s)<0$, lemma 2 ensures that ${\mathcal{A}_{S}^{D}}_{k}(s)=\mathcal{A}$.
@@
Case 2: $\phi(s)=\max_{i}\phi_{i}(s)<0$. Therefore, we have $\phi_{i}(s)<0$ for all $i$. According to Lemma 2, $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}$. Hence the set of safe control satisfies the following relationship
@@
In summary, $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$ and the claim holds. ∎
@@ C.2 Feasibility of ISSA
Corollary 2 (Feasibility of Algorithm 3).
@@
Note that Phase 2 first finds an anchor safe control $a^{u}$, then use it with AdamBA (Algorithm 2) to find the solution of (6). In the following discussion, we first show that the safety index design rule guarantees $a^{u}$ can be found with finite iterations (Lemma 3). Then we show that AdamBA will return a solution if it enters the exponential decay phase (Lemma 4). Subsequently, we show that the evoked AdamBA procedures in phase 2 will definitely enter exponential decay phase (Lemma 5). Finally, we provide a upper bound of the computation iterations for Algorithm 3.
Lemma 3 (Existence).
@@
Proof.
According to line 13-15, after an anchor safe control is being sampled, phase 2 of Algorithm 3 will evoke at most two AdamBA processes. Hence, Lemma 5 can be proved by showing one of the two AdamBA will return a local optimal solution for (6). Next we show Lemma 5 holds in two cases.
@@
In this case, the first AdamBA process finds a safe control $a^{u}$ solution (the return of AdamBA is a set, whereas the set here has at most one element). According to Algorithm 2, a solution will be returned only if AdamBA enters exponential decay stage. Hence, according to Lemma 4, $a^{u}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
@@
In this case, the second AdamBA process is evoked (line 15 of Algorithm 3). Since no solution is returned from the first AdamBA process (13 of Algorithm 3), where we start from $a^{r}$ and exponentially outreach along the direction $\vec{v}_{a}=\frac{a^{u}-a^{r}}{||a^{u}-a^{r}||}$, then all the searched control point along $\vec{v}_{a}$ is UNSAFE.
@@
During the second AdamBA process (line 15 of Algorithm 3), we start from $a^{u}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{a^{r}-a^{u}}{||a^{r}-a^{u}||}$ with initial step size $\beta=\frac{\|a^{r}-a^{u}\|}{4}$. Hence, denote $\bar{a}_{s}^{1}$ as the first sample point along $\vec{v}_{r}$, and $\bar{a}_{s}^{1}$ satisfies
@@
which indicates $\bar{a}_{s}^{1}=a_{s}^{2}$, and the safety status of $\bar{a}_{s}^{1}$ is UNSAFE. Since the safe status of $a^{u}$ is SAFE and a UNSAFE point can be sampled during the exponential outreach stage, the second AdamBA process (line 15 of Algorithm 3) will enter exponential decay stage. Therefore, according to Lemma 4, ${a^{u}}^{*}$ will always be returned and ${a^{u}}^{*}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
@@
Proof of Corollary 2.
Proof.
According to Lemma 3 and Lemma 5, Algorithm 3 is able to find local optimal solution of (6). Next, we will prove Algorithm 3 can be finished within finite iterations.
According to Algorithm 3, ISSA include (i) one procedure to find anchor safe control $a_{a}$, and (ii) at most three AdamBA procedures. Firstly, based on Lemma 3, $a_{a}$ can be found within finite iterations. Secondly, each AdamBA procedure can be finished within finite iterations due to:
@@
C.3 Proof of Theorem 1
Before we prove Theorem 1, we first define set $\mathcal{S}_{S}^{D}:=\{s|\phi(s)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{S}_{S}^{D}$ that is useful for proving the main theorem:
Lemma 6 (Forward Invariance of $\mathcal{S}_{S}^{D}$).
@@ C.3 Proof of Theorem 1
Proof.
If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, then we can ensure the system has nonempty set of safe control at any state by Corollary 1. By Corollary 2, the implicit safe set algorithm can always find local optima solution of (6). The local optima solution always satisfies the constraint $\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}$, which indicates that 1) if $\phi(s_{t_{0}})\leq 0$, then $\phi(s_{t})\leq 0,\forall t\geq t_{0}$. Note that $\phi(s)\leq 0$ demonstrates that $s\in\mathcal{S}_{S}^{D}$. ∎
Proof the Theorem 1
Proof.
Leveraging Lemma 6, we then proceed to prove that the forward invariance to the set $\mathcal{S}_{S}^{D}$ guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$ which $\mathcal{\bar{S}}=\mathcal{S}_{S}\cap\mathcal{S}_{S}^{D}$. Depending on the relationship between $\mathcal{S}_{S}^{D}$ and $\mathcal{S}_{S}$, there are two cases in the proof which we will discuss below.
@@
In this case, $\mathcal{\bar{S}}=\mathcal{S}_{S}^{D}$. According to Lemma 6, If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$ and hence $\mathcal{\bar{S}}$.
@@
Summarizing the above two cases, we have shown that if ${\phi_{0}}_{i}(s_{t})\leq 0$ then ${\phi_{0}}_{i}(s_{t+1})\leq 0$, which indicates if $\forall i,{\phi_{0}}_{i}(s_{t})\leq 0$ then $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$. Note that $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$ indicates that $\phi_{0}(s_{t+1})=\max_{i}{\phi_{0}}_{i}(s_{t+1})\leq 0$. Therefore, we have that if $s_{t}\in\mathcal{\bar{S}}$ then $s_{t+1}\in\mathcal{S}_{S}$. Thus, we also have $s_{t+1}\in\mathcal{\bar{S}}$ by Lemma 6. By induction, we have if $s_{t_{0}}\in\mathcal{\bar{S}}$, $s_{t}\in\mathcal{\bar{S}},\forall t>t_{0}$.
@@
Appendix D Proof of Theorem 2

### v4

@@ 7 Conclusion and Future Prospectus
In this study, we introduce Safe Set Guided State-wise Constrained Policy Optimization (S-3PO), a novel algorithm pioneering state-wise safe optimal policies. This distinction is underlined by the absence of training violations, signifying an error-free learning paradigm. S-3PO employs a safeguard anchored in black-box dynamics to ensure secure exploration. Subsequently, it integrates a novel “imaginary” safety cost to guide the RL agent towards optimal safe policies. In the following work, we will try to implement our algorithm in real robot implementation.
@@ A.1 Circumstances and Assumptions
Assumption 1 (2D Collision Avoidance)
@@ A.2 Safety Index Synthesis
Following the rules in (Liu and Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{A}_{S}^{D}(s)$ is always nonempty. By setting $\eta=0$, the parameterization rule of safety index design rule is defined as:
@@ B.1 S-3PO Algorithm
[…]
           $\theta_{k+1}\leftarrow\theta^{\prime}$Update policy}
@@ B.2 Additional Algorithms
1:  procedure: AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\vec{v}^{r},S,S_{goal}$)
@@
1:  procedure: ISSA($\epsilon,\beta,\Sigma,n,a^{r}$)
2:     Phase 1: {Phase 1}
@@
20:  end procudure
Appendix C Proof of Theorem 5.1
To prove the Theorem 5.1 we introduce two important corollaries to show that 1) the set of safe control is always nonempty if we choose a safety index that satisfies the design rule in (11); and 2) the proposed algorithm 3 is guaranteed to find a safe control if there exists one. With these two corollaries, it is then straightforward to prove the Theorem 5.1.
@@ C.1 Feasibility of the Safety Index for Continuous-Time System
Corollary C.1 (Non-emptiness of the set of safe control).
@@
Note that the set of safe control $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$ is non-empty if and only if it is non-empty in the following two cases: $\phi(s)-\eta<0$ or $\phi(s)-\eta\geq 0$. In the following discussion, we first show that the safety index design rule guarantees a non-empty set of safe control if there’s only one obstacle when $\phi(s)-\eta\geq 0$ (Lemma C.2). Then we show that the set of safe control is non-empty if there’s only one obstacle when $\phi(s)-\eta<0$ (Lemma C.4). Finally, we leverage Lemma C.2 and Lemma C.4 to show $\mathcal{A}_{S}^{D}(s)$ is non-empty if there’re multiple obstacles at any state.
Lemma C.2.
@@
The three cases cover all possible situations. Hence (16) always hold and the claim in the lemma is verified.
Lemma C.4.
@@
Now we show that $\dot{\phi}$ is bounded. According to the definition of safety index design rule, $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$. We ignored the subscript $i$ since it is assumed that there is only one obstacle. According to Assumption 1, we have the state space is bounded, thus both $d$ and $\dot{d}$ are bounded, which implies that $nd^{n-1}\dot{d}$ is bounded. Moreover, we have for all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$ according to Assumption 1, which indicates the mapping from $a$ to $(z,w)$ is surjective. Since $z$ and $w$ are bounded and both can achieve zeros according to Assumption 1, we have afor-$\forall a$, the corresponding $(z,w)$ are bounded. Since $\ddot{d}=-z\cos(\alpha)+v\sin(\alpha)w$, then $\ddot{d}$ is bounded. Hence $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ any $s$ that $\phi(s)-\eta<0$ and the claim is true.
Proof of Corollary C.1.
Proof C.6.
If there is one obstacle, then lemma C.2 and lemma C.4 have proved that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for all $s$. Now we need to consider the case where there are more than one obstacle but the environment is sparse in the sense that at any time step, there is at most one obstacle which is safety critical, i.e. $\phi_{i}\geq 0$. To show nonemptiness of $\mathcal{A}_{S}^{D}(s)$, we consider the following two cases. In the following discussion, we set $\eta=0$ according to the safety index design rule.
@@
Denote ${\mathcal{A}_{S}^{D}}_{j}(s):=\{a\in\mathcal{A}\mid\phi_{j}(f(s,a))\leq\phi_{j}(s)\}$. Lemma C.2 ensures that ${\mathcal{A}_{S}^{D}}_{j}(s)$ is nonempty. Denote ${\mathcal{A}_{S}^{D}}_{k}(s):=\{a\in\mathcal{A}\mid\phi_{k}(f(s,a))\leq 0\}$ where $k\neq j$. Since $\phi_{k}(s)<0$, lemma C.4 ensures that ${\mathcal{A}_{S}^{D}}_{k}(s)=\mathcal{A}$.
@@
Case 2: $\phi(s)=\max_{i}\phi_{i}(s)<0$. Therefore, we have $\phi_{i}(s)<0$ for all $i$. According to Lemma C.4, $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}$. Hence the set of safe control satisfies the following relationship
@@
In summary, $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$ and the claim holds.
@@ C.2 Feasibility of ISSA
Corollary C.7 (Feasibility of Algorithm 3).
@@
Note that Phase 2 first finds an anchor safe control $a^{u}$, then use it with AdamBA (Algorithm 2) to find the solution of (5). In the following discussion, we first show that the safety index design rule guarantees $a^{u}$ can be found with finite iterations (Lemma C.8). Then we show that AdamBA will return a solution if it enters the exponential decay phase (Lemma C.10). Subsequently, we show that the evoked AdamBA procedures in phase 2 will definitely enter exponential decay phase (Lemma C.12). Finally, we provide a upper bound of the computation iterations for Algorithm 3.
Lemma C.8 (Existence).
@@
Proof C.13.
According to line 13-16, after an anchor safe control is being sampled, phase 2 of Algorithm 3 will evoke at most two AdamBA processes. Hence, Lemma C.12 can be proved by showing one of the two AdamBA will return a local optimal solution for (5). Next we show Lemma C.12 holds in two cases.
@@
In this case, the first AdamBA process finds a safe control $a^{u}$ solution (the return of AdamBA is a set, whereas the set here has at most one element). According to Algorithm 2, a solution will be returned only if AdamBA enters exponential decay stage. Hence, according to Lemma C.10, $a^{u}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
@@
In this case, the second AdamBA process is evoked (line 16 of Algorithm 3). Since no solution is returned from the first AdamBA process (13 of Algorithm 3), where we start from $a^{r}$ and exponentially outreach along the direction $\vec{v}_{a}=\frac{a^{u}-a^{r}}{||a^{u}-a^{r}||}$, then all the searched control point along $\vec{v}_{a}$ is UNSAFE.
@@
During the second AdamBA process (line 16 of Algorithm 3), we start from $a^{u}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{a^{r}-a^{u}}{||a^{r}-a^{u}||}$ with initial step size $\beta=\frac{\|a^{r}-a^{u}\|}{4}$. Hence, denote $\bar{a}_{s}^{1}$ as the first sample point along $\vec{v}_{r}$, and $\bar{a}_{s}^{1}$ satisfies
@@
which indicates $\bar{a}_{s}^{1}=a_{s}^{2}$, and the safety status of $\bar{a}_{s}^{1}$ is UNSAFE. Since the safe status of $a^{u}$ is SAFE and a UNSAFE point can be sampled during the exponential outreach stage, the second AdamBA process (line 16 of Algorithm 3) will enter exponential decay stage. Therefore, according to Lemma C.10, ${a^{u}}^{*}$ will always be returned and ${a^{u}}^{*}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
@@
Proof of Corollary C.7.
Proof C.14.
According to Lemma C.8 and Lemma C.12, Algorithm 3 is able to find local optimal solution of (5). Next, we will prove Algorithm 3 can be finished within finite iterations.
According to Algorithm 3, ISSA include (i) one procedure to find anchor safe control $a_{a}$, and (ii) at most three AdamBA procedures. Firstly, based on Lemma C.8, $a_{a}$ can be found within finite iterations. Secondly, each AdamBA procedure can be finished within finite iterations due to:
@@
C.3 Proof of Theorem 5.1
Before we prove Theorem 5.1, we first define set $\mathcal{S}_{S}^{D}:=\{s|\phi(s)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{S}_{S}^{D}$ that is useful for proving the main theorem:
Lemma C.15 (Forward Invariance of $\mathcal{S}_{S}^{D}$).
@@
Proof C.16.
If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, then we can ensure the system has nonempty set of safe control at any state by Corollary C.1. By Corollary C.7, the implicit safe set algorithm can always find local optima solution of (5). The local optima solution always satisfies the constraint $\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}$, which indicates that 1) if $\phi(s_{t_{0}})\leq 0$, then $\phi(s_{t})\leq 0,\forall t\geq t_{0}$. Note that $\phi(s)\leq 0$ demonstrates that $s\in\mathcal{S}_{S}^{D}$.
Proof the Theorem 5.1
Proof C.17.
Leveraging Lemma C.15, we then proceed to prove that the forward invariance to the set $\mathcal{S}_{S}^{D}$ guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$ which $\mathcal{\bar{S}}=\mathcal{S}_{S}\cap\mathcal{S}_{S}^{D}$. Depending on the relationship between $\mathcal{S}_{S}^{D}$ and $\mathcal{S}_{S}$, there are two cases in the proof which we will discuss below.
@@
In this case, $\mathcal{\bar{S}}=\mathcal{S}_{S}^{D}$. According to Lemma C.15, If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$ and hence $\mathcal{\bar{S}}$.
@@
Summarizing the above two cases, we have shown that if ${\phi_{0}}_{i}(s_{t})\leq 0$ then ${\phi_{0}}_{i}(s_{t+1})\leq 0$, which indicates if $\forall i,{\phi_{0}}_{i}(s_{t})\leq 0$ then $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$. Note that $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$ indicates that $\phi_{0}(s_{t+1})=\max_{i}{\phi_{0}}_{i}(s_{t+1})\leq 0$. Therefore, we have that if $s_{t}\in\mathcal{\bar{S}}$ then $s_{t+1}\in\mathcal{S}_{S}$. Thus, we also have $s_{t+1}\in\mathcal{\bar{S}}$ by Lemma C.15. By induction, we have if $s_{t_{0}}\in\mathcal{\bar{S}}$, $s_{t}\in\mathcal{\bar{S}},\forall t>t_{0}$.
@@
Appendix D Proof of Theorem 5.2